OpenStudy (anonymous):

does (x – 3)(x^2 – 4x – 7) simplify to x^3 – 7x^2 + 19x + 21..???

OpenStudy (whpalmer4):

Well, do the multiplication and find out! \[(x-3)(x^2-4x-7) = x(x^2-4x-7) -3(x^2-4x-7)\]\[\qquad =\]